OpenStudy (anonymous):

Help please! Use the quotient property to rewrite the following algebraic expression. log4(7/3)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Use the rule \[\Large \log_{b}\left(\frac{x}{y}\right) = \log_{b}(x)-\log_{b}(y)\]

OpenStudy (anonymous):

i have never seen that im confused.

jimthompson5910 (jim_thompson5910):

it was never brought up during the lesson?

OpenStudy (anonymous):

i do online school so its harder for me to understand.. math has always been my worst subject

jimthompson5910 (jim_thompson5910):

an example of that rule in action log(12/7) = log(12) - log(7) this works for any log base

OpenStudy (anonymous):

If i dont have a calculator how would i do it?

jimthompson5910 (jim_thompson5910):

you don't need a calculator to use the rule

jimthompson5910 (jim_thompson5910):

the rule is log(x/y) = log(x) - log(y)

jimthompson5910 (jim_thompson5910):

replace x and y with the values that make up the fraction

OpenStudy (anonymous):

ohhh i get it. thank you